Suppose that the seller decrease Price for his/her good by5 percent.Discuss how total revenue changes when 1) ed=0.3 2) end =1 3). ed =1.5

1
Expert's answer
2022-03-14T13:16:38-0400

Solution

When Ed=0.3, percentage change in quantity will slightly increase by

 "0.3\times 5\%=1.5\%" ,

So total revenue will increase by 

"0.3\times 5\%=1.5\%"


when Ed=1, total revenue will increase by

 "1\times5=5\%" 

when prices decrease by the said percentage.


when Ed=1.5, total revenue will increase by

 "1.5\times 5\%=7.5\%" 

Since an elasticity of demand above 1 shows customers react more to changes in price.